# 2.5cm2.5cm2cm2cm

### v1

@@ 3.1. The Nash discriminant variety
Define the morphism $\psi\colon{\mathcal{C}}\to{\mathcal{L}}$ by $\psi(p)\coloneqq[f]$ if and only if $p\in Z(f)$. This is a finite morphism of degree $c(2,2,2)=2$. If $[f]\in\Delta(E)\cap{\mathcal{L}}$, then, by assumption, the zero scheme $Z(f)$ of $f$ is a nonreduced point of multiplicity $2$ supported at a point $p\in{\mathcal{C}}$. To put it another way, this means that the finite morphism $\psi$ is ramified at $p$ with ramification index $2$. Therefore, finding the intersection number $|\Delta(E)\cap{\mathcal{L}}|$ is equivalent to counting the number of branch points of $\psi$. Furthermore, since the ramification index of $\psi$ at each ramification point is $2$, the number of branch points of $\psi$ coincides with the degree of the ramification divisor of $\psi$. It follows from the Hurwitz-Riemann formula [Har77, Chapter IV, Section 2] that the degree of the ramification divisor of $\psi$, and hence the degree of $\Delta(E)$, is $(2-2\,g({\mathcal{L}}))\deg\psi+2\,g({\mathcal{C}})-2=6$. $\diamondsuit$
@@
It can be verified, as in Example 3.9, that the dependency locus ${\mathcal{C}}$ of two generic global sections $f_{1}$ and $f_{2}$ of $E$ is a nonsingular irreducible curve of genus $1$ in ${\mathbb{P}}^{\boldsymbol{d}}$ by calculating the Eagon-Northcott complex of the sheaf morphism $(f_{1},f_{2})\colon{\mathcal{O}}^{\oplus 2}\to E$. The map $\psi\colon{\mathcal{C}}\to{\mathcal{L}}$ defined in the same way as in Example 3.9 is a finite morphism of degree $c(2,2,3)=2$. Thus, the Hurwitz-Riemann formula shows that the degree of the ramification divisor of $\psi$, and hence the degree of $\Delta(E)$, is $4$diamondsuit
@@ 3.2. The Nash discriminant variety of a three-player binary game
In this section, we focus on three-player games of format ${\boldsymbol{d}}=(2,2,2)$. Thanks to Remark 3.4, the study of $\Delta(2,2,2)$ is equivalent to the study of $\Delta(E)$, where in this format the vector bundle $E$ on ${\mathbb{P}}^{\boldsymbol{d}}$ defined in (2.4) is $E={\mathcal{O}}(0,1,1)\oplus{\mathcal{O}}(1,0,1)\oplus{\mathcal{O}}(1,1,0)$ Throughout this section, we use $\Delta$ to denote $\Delta(E)$ for simplicity. Firstly, in Theorem 3.11, we show that $\Delta\subset{\mathbb{P}}H^{0}({\mathbb{P}}^{\boldsymbol{d}},E)$ is an irreducible hypersurface of degree $6$. Secondly, since a generic global section $f$ of $E$ is such that ${\mathcal{Z}}(f)$ is a nonreduced point of multiplicity $2$, we investigate the loci of global sections such that $\dim({\mathcal{Z}}(f))>0$. In particular, in Propositions 3.15 and 3.16, we characterize all zero schemes ${\mathcal{Z}}(f)$ of global sections containing either a one-dimensional or a two-dimensional component. Lastly, we compute the codimensions, degrees, and equations and study the irreducibility of the loci of global sections $f$ of $E$ such that ${\mathcal{Z}}(f)$ is positive-dimensional and of a specific type. With the aid of Macaulay2, we utilize these results to determine all singular strata of $\Delta$ (Table 3, [APS25]).
@@
We note a similarity in our approach to the universality theorem on Nash equilibria. The theorem proves that every real algebraic variety is isomorphic to the set of totally mixed Nash equilibria of a three-player game or an $n$-player game in which each player has two strategies. However, even for a real algebraic curve, more than two strategies per player are needed in a three-player game, or more than three players in a binary game (see [Dat03, Theorem 5, Theorem 6]). While the universality result begins with a given variety and constructs a corresponding game, our approach follows a similar line of thought but takes a slightly different focus: we fix the game and examine all possible zero schemes of global sections $f$ of $E$.
@@
for each $i\in\{2,3\}$. Thus, $[f]\in\Delta$ if and only if the following binary quadratic form in $\pi^{(3)}_{1}$ and $\pi^{(3)}_{2}$ defines a nonreduced point of multiplicity $2$ in ${\mathbb{P}}(V_{3})$:
@@
Its discriminant $\Delta(f^{(3)})$ is a quadratic form in $a^{(3)}_{11},\ldots,a^{(3)}_{22}$ whose coefficients are bihomogeneous polynomials of bidegree $(2,2)$ in $a^{(1)}_{11},\ldots,a^{(1)}_{22}$ and $a^{(2)}_{11},\ldots,a^{(2)}_{22}$. This quadratic form defines a nonsingular quadric surface in ${\mathbb{P}}H^{0}({\mathbb{P}}^{\boldsymbol{d}},{\mathcal{O}}({\boldsymbol{1}}_{3}))$ if and only if the $4\times 4$ symmetric matrix of the quadratic form has rank $4$. The complement ${\mathcal{U}}_{0}$ of the hypersurface defined by the $4\times 4$ symmetric matrix in ${\mathcal{U}}$ is a nonempty open subset (for example, the symmetric matrix has rank $4$ if $f^{(1)}=\pi^{(2)}_{1}\pi^{(3)}_{1}-\pi^{(2)}_{1}\pi^{(3)}_{2}-\pi^{(2)}_{2}\pi^{(3)}_{1}-\pi^{(2)}_{2}\pi^{(3)}_{2}$ and $f^{(2)}=-\pi^{(1)}_{1}\pi^{(3)}_{1}+\pi^{(1)}_{1}\pi^{(3)}_{2}+\pi^{(1)}_{2}\pi^{(3)}_{1}+\pi^{(1)}_{2}\pi^{(3)}_{2}$).
The fiber of $\Psi$ over any point $[(f^{(1)},f^{(2)})]$ of ${\mathcal{U}}_{0}$,
@@
One verifies by direct computations that ${\mathcal{Z}}(f)$ is a nonreduced point of multiplicity $2$ supported at $[{\boldsymbol{\pi}}]=([0:1],[1:-1]:[1:3])\in{\mathbb{P}}^{\boldsymbol{d}}$. This can be verified by observing that the determinant of the matrix obtained evaluating (3.7) at the coefficients of $f$ is zero. This implies $[f]\in\Delta$. The point $[{\boldsymbol{\pi}}]$ cannot be associated with a triple of probability distributions. Hence, the game $X$ does not admit totally mixed Nash equilibria. However, it has mixed Nash equilibria, which is the union of two line segments containing two pure Nash equilibria. This makes us ask the following natural question. $\diamondsuit$
@@
gives rise to $\dim H^{0}({\mathcal{C}},{\mathcal{O}}_{\mathcal{C}}(1))=4$ and implies that the Hilbert polynomial of ${\mathcal{C}}$ is $13t+$ 1. Therefore, it follows that ${\mathcal{C}}$ is a space curve of dimension $1$, degree $3$, and arithmetic genus 00, and hence it is a (possibly degenerate) twisted cubic in the Segre embedding.
@@
By assumption, $H^{0}({\mathbb{P}}^{\boldsymbol{d}},E(-{\boldsymbol{a}}))\not=0$. Thus, either (1) two of the integers in ${\boldsymbol{a}}$ are $1$, and the remaining one is 00, or (2) one of the integers in ${\boldsymbol{a}}$ is $1$, and the remaining two are both 00.
@@
Type $\star$ $\dim\Delta^{\star}${components} $\deg\Delta^{\star}$ description of ${\mathcal{Z}}(f)$, $f$ generic on $\Delta^{\star}$
$\mathrm{cub}$ $8$ $1$ $11^{*}$ Nonsingular twisted cubic, Proposition 3.18
$\mathrm{con}$ $8$ $3$ $12=4\times 3$ Nonsingular conic, Proposition 3.20
$\mathrm{lin}$ $8$ $3$ $24=8\times 3$ Line, Proposition 3.22
$\mathrm{ll}$ $7$ $3$ $24=8\times 3$ Singular plane conic, [APS25, Proposition 1]
$\mathrm{cl}$ $7$ $3$ $30=10^{*}\times 3$ Space connected cubic, union of a nonsingular conic and a line, [APS25, Proposition 2]
$\mathrm{lll}$ $6$ $3$ $30=10^{*}\times 3$ Space connected cubic, union of three nonconcurrent lines, [APS25, Proposition 3]
$\mathrm{3l}$ $5$ $1$ $14^{*}$ Union of three nonplanar, concurrent lines, [APS25, Proposition 4]
$\mathrm{ql}$ $4$ $3$ $12=4\times 3$ Union of a quadric surface and a line,
@@
$\mathrm{scr}$ $3$ $3$ $3=1\times 3$ Rational normal scroll, Proposition 3.17(1)
$\mathrm{qq}$ $2$ $3$ $6=2\times 3$ Union of two nonsingular quadrics,
@@
$\mathrm{scr}$ Rational normal scroll/union of two quadric surfaces $2$
$\mathrm{ql}$ Quadric surface $3$
$\mathrm{cub}$ Possibly degenerate twisted cubic $4$
$\mathrm{con}$ Possibly degenerate plane conic $5$
$\mathrm{lin}$ Line $6$
@@
We prove that $\dim\,{\mathcal{U}}=8$. Define the rational map $\Psi\colon{\mathbb{P}}H^{0}({\mathbb{P}}^{\boldsymbol{d}},E)\dashrightarrow{\mathbb{P}}^{7}={\mathbb{P}}((V_{1}^{*}\oplus V_{3}^{*})\otimes(V_{1}^{*}\oplus V_{2}^{*}))$ by $\Psi([f])=[(f^{(2)},f^{(3)})]$ if $f=(f^{(1)},f^{(2)},f^{(3)})$. The image ${\mathcal{U}}$ under $\Psi$, denoted ${\mathcal{U}}^{\prime}$, is an open subset of ${\mathbb{P}}^{7}$, and hence it has dimension $7$. Thus, it suffices to show that for any point $[f]\in{\mathcal{U}}$, the fiber of the morphism obtained from $\Psi$ by restricting to ${\mathcal{U}}$ over $\Psi([f])$ has dimension $1$.
@@
Suppose that $[f^{\prime}]\in\Psi^{-1}(\Psi([f]))$. Let $(f^{\prime})^{(1)}$ be the first coordinate function of $f^{\prime}$. The diagonal embeddings of the lines in $\Pi$ associated with $f$ and $f^{\prime}$ intersect nontrivially, so they must share elements. Thus, there exist linear forms $L$ and $L^{\prime}$ in $\pi^{(1)}$ such that $Lf^{(1)}=L^{\prime}(f^{\prime})^{(1)}$, from which it follows that $f^{(1)}$ and $(f^{\prime})^{(1)}$ (and hence $L$ and $L^{\prime}$) differ only by a scalar multiple. Therefore, we have
@@
For each $i\in[3]$, denote by $\Delta_{i}^{\mathrm{con}}$ the irreducible component of $\Delta^{\mathrm{con}}$ which contains a global section $f=(f^{(1)},f^{(2)},f^{(3)})$ of $E$ such that $f^{(i)}$ is irreducible, but the others are reducible and share a linear factor. This remark concerns a determinantal expression for the equations for $\Delta_{i}^{\mathrm{con}}$. We only discuss the case when $i=1$ as the remaining cases can be treated similarly.
@@
The element $[(f^{(3)},f^{(2)})]\in{\mathbb{P}}(V_{1}^{*}\otimes(V_{2}^{*}\oplus V_{3}^{*}))$ is contained in the image of the Segre map if and only if $(f^{(3)},f^{(2)})$ has rank $1$ when considering it as a linear transformation from $V_{1}$ to $V_{3}^{*}\oplus V_{2}^{*}$. The latter is equivalent to the condition that the $2\times 2$ minors of the matrix representation of the linear transformation relative to given bases for $V_{1}$ and $V_{3}^{*}\oplus V_{2}^{*}$ are all zero. More specifically, if $f^{(2)}=\sum_{j,k=0}^{1}a_{jk}^{(2)}\pi^{(1)}_{j}\pi^{(3)}_{k}$ and let $f^{(3)}=\sum_{j,k=0}^{1}a_{jk}^{(3)}\pi^{(1)}_{j}\pi^{(2)}_{k}$, then the matrix of the linear transformation $(f^{(3)},f^{(2)})\colon V_{1}\to V_{3}^{*}\oplus V_{2}^{*}$ relative to $\{e^{(1)}_{1},e^{(1)}_{2}\}$ and $\{\pi^{(3)},\pi^{(2)}\}$ is
@@
Let ${\mathcal{U}}$ be the open subset of points of $\Delta_{1}^{\mathrm{lin}}$ of the form $[(MN^{\prime}+M^{\prime}N,LN,L^{\prime}M)]$ with $L,L^{\prime}\in V_{1}^{*}\setminus\{{\boldsymbol{0}}\},M,M^{\prime}\in V_{2}^{*}\setminus\{{\boldsymbol{0}}\}$, and $N,N^{\prime}\in V_{3}^{*}\setminus\{{\boldsymbol{0}}\}$, where both $\{M,M^{\prime}\}$ and $\{N,N^{\prime}\}$ are linearly independent so that $MN^{\prime}+M^{\prime}N$ is irreducible. Define the map $\Psi\colon{\mathcal{U}}\to{\mathbb{P}}((V_{1}^{*}\otimes V_{3}^{*})\oplus(V_{1}^{*}\otimes V_{2}^{*}))$ by $\Psi([(MN^{\prime}+M^{\prime}N,LN,L^{\prime}M)])=[(LN,L^{\prime}M)]$. The image of this map is the complete intersection of two quadric hypersurfaces; one is the cone over the Segre embedding of ${\mathbb{P}}(V_{1}^{*})\times{\mathbb{P}}(V_{3}^{*})$ in ${\mathbb{P}}(V_{1}^{*}\otimes V_{3}^{*})$ given by sending $([L],[N])$ to $[LN]$, and the other is the cone over the Segre embedding of ${\mathbb{P}}(V_{1}^{*})\times{\mathbb{P}}(V_{2}^{*})$ in ${\mathbb{P}}(V_{1}^{*}\otimes V_{2}^{*})$ by sending $([L^{\prime}],[M])$ to $[L^{\prime}M]$. So, it is irreducible of dimension $5$. The fiber of $\Psi$ over $[(LN,L^{\prime}M)]$ is
@@
This remark concerns the equations for $\Delta_{i}^{\mathrm{lin}}$ for each $i\in[3]$. We only discuss the case when $i=1$ as the remaining cases can similarly be done. With the aim of doing so, we use (3.6) to describe the algebraic relations among the coefficients of the components of $f=(f^{(1)},f^{(2)},f^{(3)})\in H^{0}({\mathbb{P}}^{\boldsymbol{d}},E)$ such that ${\mathcal{Z}}(f)$ is a line.
@@
The condition that $f^{(2)}$ factors into two linear forms is the same as the condition that $[f^{(2)}]$ lies in the Segre embedding of ${\mathbb{P}}(V_{1}^{*})\times{\mathbb{P}}(V_{3}^{*})$ in ${\mathbb{P}}(V_{1}^{*}\otimes V_{3}^{*})$. The latter is equivalent to the condition that $f^{(2)}$ has rank one when considering it as a linear transformation from $V_{1}$ to $V_{3}^{*}$. Thus, the condition for $f^{(2)}$ to be a product of linear forms is expressed as the vanishing of the determinant of the matrix of this linear transformation relative to the basis $\{e^{(1)}_{1},e^{(1)}_{2}\}$ for $V_{1}$ and the basis $\{\pi^{(3)}_{1},\pi^{(3)}_{2}\}$ for $V_{3}^{*}$:
@@
The first coordinate function $f^{(1)}$ can be expressed as a linear combination of $M$ and $N$ if and only if $f^{(1)}(p_{M},p_{N})=0$, where $p_{M}\in V_{2}$ and $p_{N}\in V_{3}$ such that $[p_{M}]$ is the point of ${\mathbb{P}}(V_{2})$ defined by $M$ and $[p_{N}]$ the point of ${\mathbb{P}}(V_{3})$ defined by $N$. The latter can be translated into the system of polynomial equations
$\begin{cases}F_{11}\coloneqq f^{(1)}\left(a^{(2)}_{12},-a^{(2)}_{11},a^{(3)}_{12},-a^{(3)}_{11}\right)=0\\[4.0pt] F_{12}\coloneqq f^{(1)}\left(a^{(2)}_{12},-a^{(2)}_{11},a^{(3)}_{22},-a^{(3)}_{21}\right)=0\\[4.0pt] F_{21}\coloneqq f^{(1)}\left(a^{(2)}_{22},-a^{(2)}_{21},a^{(3)}_{12},-a^{(3)}_{11}\right)=0\\[4.0pt] F_{22}\coloneqq f^{(1)}\left(a^{(2)}_{22},-a^{(2)}_{21},a^{(3)}_{22},-a^{(3)}_{21}\right)=0\,,\end{cases}$
@@
$[N]=\begin{cases}\left[a^{(2)}_{11}\pi^{(2)}_{1}+a^{(2)}_{12}\pi^{(2)}_{2}\right]&\text{if $\left(a^{(2)}_{11},a^{(2)}_{12}\right)\neq(0,0)$}\,,\\[4.0pt] \left[a^{(2)}_{21}\pi^{(2)}_{1}+a^{(2)}_{22}\pi^{(2)}_{2}\right]&\text{if $\left(a^{(2)}_{21},a^{(2)}_{22}\right)\neq(0,0)$,}\end{cases}$
@@
$[M]=\begin{cases}\left[a^{(3)}_{11}\pi^{(3)}_{1}+a^{(3)}_{12}\pi^{(3)}_{3}\right]&\text{if $\left(a^{(3)}_{11},a^{(3)}_{12}\right)\neq(0,0)$}\,,\\[4.0pt] \left[a^{(3)}_{21}\pi^{(3)}_{1}+a^{(3)}_{22}\pi^{(3)}_{2}\right]&\text{if $\left(a^{(3)}_{21},a^{(3)}_{22}\right)\neq(0,0)$.}\end{cases}$

### v2

@@ 3.1. The Nash discriminant variety
Define the morphism $\psi\colon{\mathcal{C}}\to{\mathcal{L}}$ by $\psi(p)\coloneqq[f]$ if and only if $p\in Z(f)$. This is a finite morphism of degree $c(2,2,2)=2$. If $f]\in\Delta(E)\cap{\mathcal{L}}$, then, by assumption, the zero scheme $Z(f)$ of $f$ is a nonreduced point of multiplicity $2$ supported at a point $p\in{\mathcal{C}}$. To put it another way, this means that the finite morphism $\psi$ is ramified at $p$ with ramification index $2$. Therefore, finding the intersection number $|\Delta(E)\cap{\mathcal{L}}|$ is equivalent to counting the number of branch points of $\psi$. Furthermore, since the ramification index of $\psi$ at each ramification point is $2$, the number of branch points of $\psi$ coincides with the degree of the ramification divisor of $\psi$. It follows from the Hurwitz-Riemann formula [Har77, Chapter IV, Section 2] that the degree of the ramification divisor of $\psi$, and hence the degree of $\Delta(E)$, is $(2-2\,g({\mathcal{L}}))\deg\psi+2\,g({\mathcal{C}})-2=6$. $\diamondsuit$
@@
It can be verified, as in Example 3.9, that the dependency locus $\mathcal{C}$} of two generic global sections $f_{1}$ and $f_{2}$ of $E$ is a nonsingular irreducible curve of genus $1$ in ${\mathbb{P}}^{\boldsymbol{d}}$ by calculating the Eagon-Northcott complex of the sheaf morphism $(f_{1},f_{2})\colon{\mathcal{O}}^{\oplus 2}\to E$. The map $\psi\colon{\mathcal{C}}\to{\mathcal{L}}$ defined in the same way as in Example 3.9 is a finite morphism of degree $c(2,2,3)=2$. Thus, the Hurwitz-Riemann formula shows that the degree of the ramification divisor of $\psi$, and hence the degree of $\Delta(E)$, is $4$. $\diamondsuit$
@@ 3.2. The Nash discriminant variety of a three-player binary game
In this section, we focus on three-player games of format ${\boldsymbol{d}}=(2,2,2)$. Thanks to Remark 3.4, the study of $\Delta(2,2,2)$ is equivalent to the study of $\Delta(E)$, where in this format the vector bundle $E$ on ${\mathbb{P}}^{\boldsymbol{d}}$ defined in (2.4) is $E={\mathcal{O}}(0,1,1)\oplus{\mathcal{O}}(1,0,1)\oplus{\mathcal{O}}(1,1,0)$ . Throughout this section, we use $\Delta$ to denote $\Delta(E)$ for simplicity. Firstly, in Theorem 3.11, we show that $\Delta\subset{\mathbb{P}}H^{0}({\mathbb{P}}^{\boldsymbol{d}},E)$ is an irreducible hypersurface of degree $6$. Secondly, since a generic global section $f$ of $E$ is such that ${\mathcal{Z}}(f)$ is a nonreduced point of multiplicity $2$, we investigate the loci of global sections such that $\dim({\mathcal{Z}}(f))>0$. In particular, in Propositions 3.15 and 3.16, we characterize all zero schemes ${\mathcal{Z}}(f)$ of global sections containing either a one-dimensional or a two-dimensional component. Lastly, we compute the codimensions, degrees, and equations and study the irreducibility of the loci of global sections $f$ of $E$ such that ${\mathcal{Z}}(f)$ is positive-dimensional and of a specific type. With the aid of Macaulay2, we utilize these results to determine all singular strata of $\Delta$ (Table 4, [APS25]).
@@
We note a similarity between our approach and the universality theorem on Nash equilibria. The theorem proves that every real algebraic variety is isomorphic to the set of totally mixed Nash equilibria of a three-player game or an $n$-player game in which each player has two strategies. However, even for a real algebraic curve, more than two strategies per player are needed in a three-player game, or more than three players in a binary game (see [Dat03, Theorem 5, Theorem 6]). While the universality result begins with a given variety and constructs a corresponding game, our approach follows a similar line of thought but takes a slightly different focus: we fix the game and examine all possible zero schemes of global sections $f$ of $E$.
@@
for each $i\in\{2,3\}$. Thus, $f]\in\Delta$ if and only if the following binary quadratic form in $\pi^{(3)}_{1}$ and $\pi^{(3)}_{2}$ defines a nonreduced point of multiplicity $2$ in ${\mathbb{P}}(V_{3})$:
@@
[…]
The fiber of $\Psi$ over any point $(f^{(1)},f^{(2)})]$ of ${\mathcal{U}}_{0}$,
@@
One verifies by direct computations that ${\mathcal{Z}}(f)$ is a nonreduced point of multiplicity $2$ supported at ${\boldsymbol{\pi}}]=([0:1],[1:-1]:[1:3])\in{\mathbb{P}}^{\boldsymbol{d}}$. This can be verified by observing that the determinant of the matrix obtained by evaluating (3.7) at the coefficients of $f$ is zero. This implies $f]\in\Delta$. The point $\boldsymbol{\pi}$}] cannot be associated with a triple of probability distributions. Hence, the game $X$ does not admit totally mixed Nash equilibria. However, it has mixed Nash equilibria, which is the union of two line segments containing two pure Nash equilibria. This prompts the following natural question. $\diamondsuit$
@@
gives rise to $\dim H^{0}({\mathcal{C}},{\mathcal{O}}_{\mathcal{C}}(1))=4$ and implies that the Hilbert polynomial of $\mathcal{C}$} is $3t+1$. Therefore, it follows that $\mathcal{C}$} is a space curve of dimension $1$, degree $3$, and arithmetic genus 0, and hence it is a (possibly degenerate) twisted cubic in the Segre embedding.
@@
By assumption, $H^{0}({\mathbb{P}}^{\boldsymbol{d}},E(-{\boldsymbol{a}}))\not=0$. Thus, either (1) two of the integers in $\boldsymbol{a}$} are $1$, and the remaining one is 0, or (2) one of the integers in $\boldsymbol{a}$} is $1$, and the remaining two are both 0.
@@
Type $\star$ $\dim\Delta^{\star}$ #{components} $\deg\Delta^{\star}$ description of ${\mathcal{Z}}(f)$, $f$ generic on $\Delta^{\star}$
$\mathrm{cub}$ 88 11 $11^{*}$ Nonsingular twisted cubic, Proposition 3.18
$\mathrm{con}$ 88 33 $12=4\times 3$ Nonsingular conic, Proposition 3.20
$\mathrm{lin}$ 88 33 $24=8\times 3$ Line, Proposition 3.22
$\mathrm{ll}$ 77 33 $24=8\times 3$ Singular plane conic, [APS25, Proposition 1]
$\mathrm{cl}$ 77 33 $30=10^{*}\times 3$ Space connected cubic, union of a nonsingular conic and a line, [APS25, Proposition 2]
$\mathrm{lll}$ 66 33 $30=10^{*}\times 3$ Space connected cubic, union of three nonconcurrent lines, [APS25, Proposition 3]
$\mathrm{3l}$ 55 11 $14^{*}$ Union of three nonplanar, concurrent lines, [APS25, Proposition 4]
$\mathrm{ql}$ 44 33 $12=4\times 3$ Union of a quadric surface and a line,
@@
$\mathrm{scr}$ 33 33 $3=1\times 3$ Rational normal scroll, Proposition 3.17(1)
$\mathrm{qq}$ 22 33 $6=2\times 3$ Union of two nonsingular quadrics,
@@
$\mathrm{scr}$ Rational normal scroll/union of two quadric surfaces 22
$\mathrm{ql}$ Quadric surface 33
$\mathrm{cub}$ Possibly degenerate twisted cubic 44
$\mathrm{con}$ Possibly degenerate plane conic 55
$\mathrm{lin}$ Line 66
@@
We prove that $\dim\,{\mathcal{U}}=8$. Define the rational map $\Psi\colon{\mathbb{P}}H^{0}({\mathbb{P}}^{\boldsymbol{d}},E)\dashrightarrow{\mathbb{P}}^{7}={\mathbb{P}}((V_{1}^{*}\oplus V_{3}^{*})\otimes(V_{1}^{*}\oplus V_{2}^{*}))$ by $\Psi([f])=[(f^{(2)},f^{(3)})]$ if $f=(f^{(1)},f^{(2)},f^{(3)})$. The image $\mathcal{U}$} under $\Psi$, denoted $\mathcal{U}$$^{\prime}$, is an open subset of ${\mathbb{P}}^{7}$, and hence it has dimension $7$. Thus, it suffices to show that for any point $f]\in{\mathcal{U}}$, the fiber of the morphism obtained from $\Psi$ by restricting to $\mathcal{U}$} over $\Psi([f])$ has dimension $1$.
@@
Suppose that $f^{\prime}]\in\Psi^{-1}(\Psi([f]))$. Let $(f^{\prime})^{(1)}$ be the first coordinate function of $f^{\prime}$. The diagonal embeddings of the lines in $\Pi$ associated with $f$ and $f^{\prime}$ intersect nontrivially, so they must share elements. Thus, there exist linear forms $L$ and $L^{\prime}$ in $\pi^{(1)}$ such that $Lf^{(1)}=L^{\prime}(f^{\prime})^{(1)}$, from which it follows that $f^{(1)}$ and $(f^{\prime})^{(1)}$ (and hence $L$ and $L^{\prime}$) differ only by a scalar multiple. Therefore, we have
@@
For each $i\in[3]$, denote by $\Delta_{i}^{\mathrm{con}}$ the irreducible component of $\Delta^{\mathrm{con}}$ which contains a global section $f=(f^{(1)},f^{(2)},f^{(3)})$ of $E$ such that $f^{(i)}$ is irreducible, but the others are reducible and share a linear factor. This remark concerns a determinantal expression for the equations for $\Delta_{i}^{\mathrm{con}}$. We discuss only the case where $i=1$ as the remaining cases can be treated similarly.
@@
The element $(f^{(3)},f^{(2)})]\in{\mathbb{P}}(V_{1}^{*}\otimes(V_{2}^{*}\oplus V_{3}^{*}))$ is contained in the image of the Segre map if and only if $(f^{(3)},f^{(2)})$ has rank $1$ when considering it as a linear transformation from $V_{1}$ to $V_{3}^{*}\oplus V_{2}^{*}$. The latter is equivalent to the condition that the $2\times 2$ minors of the matrix representation of the linear transformation relative to given bases for $V_{1}$ and $V_{3}^{*}\oplus V_{2}^{*}$ are all zero. More specifically, if $f^{(2)}=\sum_{j,k=0}^{1}a_{jk}^{(2)}\pi^{(1)}_{j}\pi^{(3)}_{k}$ and let $f^{(3)}=\sum_{j,k=0}^{1}a_{jk}^{(3)}\pi^{(1)}_{j}\pi^{(2)}_{k}$, then the matrix of the linear transformation $(f^{(3)},f^{(2)})\colon V_{1}\to V_{3}^{*}\oplus V_{2}^{*}$ relative to $\{e^{(1)}_{1},e^{(1)}_{2}\}$ and $\{\pi^{(3)},\pi^{(2)}\}$ is
@@
Let $\mathcal{U}$} be the open subset of points of $\Delta_{1}^{\mathrm{lin}}$ of the form $(MN^{\prime}+M^{\prime}N,LN,L^{\prime}M)]$ with $L,L^{\prime}\in V_{1}^{*}\setminus\{{\boldsymbol{0}}\},M,M^{\prime}\in V_{2}^{*}\setminus\{{\boldsymbol{0}}\}$, and $N,N^{\prime}\in V_{3}^{*}\setminus\{{\boldsymbol{0}}\}$, where both $\{M,M^{\prime}\}$ and $\{N,N^{\prime}\}$ are linearly independent so that $MN^{\prime}+M^{\prime}N$ is irreducible. Define the map $\Psi\colon{\mathcal{U}}\to{\mathbb{P}}((V_{1}^{*}\otimes V_{3}^{*})\oplus(V_{1}^{*}\otimes V_{2}^{*}))$ by $\Psi([(MN^{\prime}+M^{\prime}N,LN,L^{\prime}M)])=[(LN,L^{\prime}M)]$. The image of this map is the complete intersection of two quadric hypersurfaces; one is the cone over the Segre embedding of ${\mathbb{P}}(V_{1}^{*})\times{\mathbb{P}}(V_{3}^{*})$ in ${\mathbb{P}}(V_{1}^{*}\otimes V_{3}^{*})$ given by sending $([L],[N])$ to $[LN]$, and the other is the cone over the Segre embedding of ${\mathbb{P}}(V_{1}^{*})\times{\mathbb{P}}(V_{2}^{*})$ in ${\mathbb{P}}(V_{1}^{*}\otimes V_{2}^{*})$ by sending $([L^{\prime}],[M])$ to $L^{\prime}M]$. So, it is irreducible of dimension $5$. The fiber of $\Psi$ over $(LN,L^{\prime}M)]$ is
@@
This remark concerns the equations for $\Delta_{i}^{\mathrm{lin}}$ for each $i\in[3]$. We discuss only the case where $i=1$ as the remaining cases can similarly be done. With the aim of doing so, we use (3.6) to describe the algebraic relations among the coefficients of the components of $f=(f^{(1)},f^{(2)},f^{(3)})\in H^{0}({\mathbb{P}}^{\boldsymbol{d}},E)$ such that ${\mathcal{Z}}(f)$ is a line.
@@
The condition that $f^{(2)}$ factors into two linear forms is the same as the condition that $f^{(2)}]$ lies in the Segre embedding of ${\mathbb{P}}(V_{1}^{*})\times{\mathbb{P}}(V_{3}^{*})$ in ${\mathbb{P}}(V_{1}^{*}\otimes V_{3}^{*})$. The latter is equivalent to the condition that $f^{(2)}$ has rank one when considering it as a linear transformation from $V_{1}$ to $V_{3}^{*}$. Thus, the condition for $f^{(2)}$ to be a product of linear forms is expressed as the vanishing of the determinant of the matrix of this linear transformation relative to the basis $\{e^{(1)}_{1},e^{(1)}_{2}\}$ for $V_{1}$ and the basis $\{\pi^{(3)}_{1},\pi^{(3)}_{2}\}$ for $V_{3}^{*}$:
@@
The first coordinate function $f^{(1)}$ can be expressed as a linear combination of $M$ and $N$ if and only if $f^{(1)}(p_{M},p_{N})=0$, where $p_{M}\in V_{2}$ and $p_{N}\in V_{3}$ such that $p_{M}]$ is the point of ${\mathbb{P}}(V_{2})$ defined by $M$ and $p_{N}]$ the point of ${\mathbb{P}}(V_{3})$ defined by $N$. The latter can be translated into the system of polynomial equations
[…]
@@
$N]=\begin{cases}\left[a^{(2)}_{11}\pi^{(2)}_{1}+a^{(2)}_{12}\pi^{(2)}_{2}\right]&\text{if $\left(a^{(2)}_{11},a^{(2)}_{12}\right)\neq(0,0)$}\,,\\[4.0pt] \left[a^{(2)}_{21}\pi^{(2)}_{1}+a^{(2)}_{22}\pi^{(2)}_{2}\right]&\text{if $\left(a^{(2)}_{21},a^{(2)}_{22}\right)\neq(0,0)$,}\end{cases}$
@@
$M]=\begin{cases}\left[a^{(3)}_{11}\pi^{(3)}_{1}+a^{(3)}_{12}\pi^{(3)}_{3}\right]&\text{if $\left(a^{(3)}_{11},a^{(3)}_{12}\right)\neq(0,0)$}\,,\\[4.0pt] \left[a^{(3)}_{21}\pi^{(3)}_{1}+a^{(3)}_{22}\pi^{(3)}_{2}\right]&\text{if $\left(a^{(3)}_{21},a^{(3)}_{22}\right)\neq(0,0)$.}\end{cases}$